. Retail price of fuel in Fiji is uniformly distributed between $2.00 and $3.50 per litre.
a. Calculate the retail average price of fuel in Fiji. (1 mark)
b. What is the probability that the price of fuel will be less than $3. Show working. (2 marks)
c. What is the probability that the price of fuel will less than $2. Show working. (2 marks)
1
Expert's answer
2020-10-08T07:43:18-0400

a) Retail average price ="\frac{(a+b)} {2}"


="\frac{(2.0+3.5)}{2}" ="\frac{5.5}{2}"


Retail average price =$2.75


b) Total number of outcomes is ;

"(3.50-2.00)=1.50"


"3-2.00=1"


Thus the probability will be = "\frac{1}{1.50}"


c) Probability will be zero since


"2-2.00=0"


Giving the probability as "\frac{0}{1.50}=0"